# FUSION: Forecast-Embedded Agent Scheduling with Service Incentive Optimization over Distributed Air-Ground Edge Networks

### v1

@@ III-B1 Modeling of SDs
Recall previous discussions, we model HuUs and MaUs respectively.
(i) For HuUs: Each HuU $u_{m}^{\mathrm{H}}\in\bm{\mathcal{U}}^{\mathrm{H}}$ is modeled by a 5-tuple $u_{m}^{\mathrm{H}}=\{e_{m}^{\mathrm{H,tx}},\,r_{m}^{\mathrm{H}},\,d_{m}^{\mathrm{H}},\,f_{m}^{\mathrm{H,loc}},\,\bm{C}_{m}^{\mathrm{H}}\}$, where $e_{m}^{\mathrm{H,tx}}$ is the uplink transmit power (Watts); $r_{m}^{\mathrm{H}}$ is the required computing workload (CPU cycles); $d_{m}^{\mathrm{H}}$ is the input data size (bits); $f_{m}^{\mathrm{H,loc}}$ is the local CPU frequency; and $\bm{C}_{m}^{\mathrm{H}}$ represents the set of reachable ESs within the communication coverage of $u_{m}^{\mathrm{H}}$.
(ii) For MaUs: Each MaU $u_{n}^{\mathrm{A}}\in\bm{\mathcal{U}}^{\mathrm{A}}$ is represented by a 6-tuple $u_{n}^{\mathrm{A}}=\{\tau_{n}^{\mathrm{ddl}},\,e_{n}^{\mathrm{A,tx}},\,r_{n}^{\mathrm{A}},\,d_{n}^{\mathrm{A}},\,f_{n}^{\mathrm{A,loc}},\,\bm{C}_{n}^{\mathrm{A}}\}$, where $\tau_{n}^{\mathrm{ddl}}$ is the task deadline time; $e_{n}^{\mathrm{A,tx}}$ is the transmit power; $r_{n}^{\mathrm{A}}$ and $d_{n}^{\mathrm{A}}$ denote the required workload and data size, respectively. $f_{n}^{\mathrm{A,loc}}$ is its local CPU capability, and $\bm{C}_{n}^{\mathrm{A}}$ is the set of SPs that can serve $u_{n}^{\mathrm{A}}$.
Note that one unique aspect appears in our paper: HuUs exhibit strong latency sensitivity, e.g., even minor delays can noticeably degrade their experience and affect their emotions. In contrast, MaUs are predominantly goal-driven and primarily require that their tasks be completed before predefined deadlines. This fundamental distinction results in heterogeneous scheduling priorities and markedly different utility structures.
@@ III-B2 Modeling of SPs
Regarding two types of SPs, we have detailed models below.
@@
(i) For ESs: Each ES $s_{i}\in\bm{\mathcal{S}}$ is characterized by a 4-tuple: $s_{i}=\{f_{i}^{\mathrm{E}},\,e_{i}^{\mathrm{E}},\,G_{i},\,l_{i}^{\mathrm{E}}\}$, where $f_{i}^{\mathrm{E}}$ is the computing capacity (e.g., CPU cycles/s); $e_{i}^{\mathrm{E}}$ is its local power consumption (Watts); In addition, each ES $s_{i}$ has a set of orthogonal subcarriers that can serve at most $G_{i}$ SDs simultaneously; and $l_{i}^{\mathrm{E}}=(x_{i}^{\mathrm{E}},y_{i}^{\mathrm{E}},h_{i}^{\mathrm{E}})$ denotes its fixed 3D location.
@@
(ii) For AgPs: each AgP $v_{k}\in\bm{\mathcal{V}}$ is modeled by a 7-tuple: $v_{k}=\{C_{k}^{\mathrm{V}},\,f_{k}^{\mathrm{V}},\,e_{k}^{\mathrm{comp}},\,e_{k}^{\mathrm{move}},\,e_{k}^{\mathrm{fly}},\,G_{k}^{\mathrm{V}},\,l_{k}^{\mathrm{V}}\}$, where $C_{k}^{\mathrm{V}}$ denotes the maximum number of UAVs carried by the vehicle; $f_{k}^{\mathrm{V}}$: computing capability of each mounted UAV; $e_{k}^{\mathrm{move}}$, $e_{k}^{\mathrm{comp}}$, and $e_{k}^{\mathrm{fly}}$ refer to power consumed for ground movement, UAV computation, and UAV flight, respectively; $G_{k}^{\mathrm{V}}$ denotes the maximum number of SDs that each UAV of $v_{k}$ can serve, owing to the use of orthogonal subcarriers, and $l_{k}^{\mathrm{V}}=(x_{k}^{\mathrm{V}},y_{k}^{\mathrm{V}},H_{0}^{\mathrm{V}})$ is the ground location of the dispatching vehicle.
@@
Modeling of offline contracts: In the offline stage, a trading contract is established between each ES $s_{i}$ and AgP $v_{k}$, denoted as $\mathbb{C}^{\mathrm{E}\leftrightarrow\mathrm{V}}_{i,k}=\{N^{\mathrm{Trad}}_{i,k},p_{i}^{\mathrm{ES}},r_{i,k}^{\mathrm{ES}}\}$, where $N^{\mathrm{Trad}}_{i,k}$ represents the quantity of traded resources, $p_{i,k}^{\mathrm{ES}}$ indicates unit payment (per task) from $s_{i}$ to contractual AgP $v_{k}$, and $r_{i,k}^{\mathrm{ES}}$ refers to the unit reward (per task) of $v_{k}$ for offering services.
IV Offline Stage: Forecasting-Driven Contracting and AgP Deployment
@@
This section systematically transforms uncertain spatio-temporal service demand into proactive, incentive-compatible, and cost-aware agent deployment decisions. Our major efforts include: (i) a predictive workload forecasting module that estimates the spatiotemporal dynamics of demands across distributed ESs; (ii) a cost-aware vehicle routing strategy based on eACO, which constructs time-window–feasible candidate associations between AgPs and ESs; and (iii) a contract-design module that leverages an incentive-compatible auction mechanism to determine long-term cooperative agreements, including winning trading pairs, allocated resources, and corresponding payment rules.
@@
To facilitate analysis, let $x^{\mathrm{Off}}_{i,k}$ denote the assignment between ESs and AgPs, where $x^{\mathrm{Off}}_{i,k}=1$ indicates that ES $s_{i}$ successfully secures the service of AgP $v_{k}$ and thereby signs an offline contract, and $x^{\mathrm{Off}}_{i,k}=0$ otherwise. For notational simplicity, we define $\bm{X}^{\mathrm{Off}}=\left\{x^{\mathrm{Off}}_{i,k}\mid s_{i}\in\bm{\mathcal{S}},v_{k}\in\bm{\mathcal{V}}\right\}$ as the assignment profile.
IV-A Utility of ESs, AgP, and Auctioneer
@@ IV-A1 Utility of ESs
In the offline stage, since delivering edge services to users and leasing computing capacity can generate both costs and revenues for ESs, we model the utility of each ES $s_{i}$ as its net profit [^3] In the online stage considered here, SDs can offload tasks directly to UAVs for execution; therefore, the ES’s utility does not account for any energy cost of transmitting tasks from the ES to UAVs., i.e., revenue from serving users minus payments for buying services from AgP $v_{k}$. Formally, the profit of ES $s_{i}$ is given by
$U^{\mathrm{ES,Off}}_{i}=\sum_{v_{k}\in\bm{\mathcal{V}}}x^{\mathrm{Off}}_{i,k}N_{i,k}^{\mathrm{Trad}}\left(\overline{p}_{i}-p^{\mathrm{ES}}_{i,k}\right)$ (1)
@@
IV-A2 Utility of AgPs
For each AgP, we model its utility in terms of four key components: (i) Vehicular moving cost: the energy cost incurred by the vehicle traveling to ES $s_{i}$ is $c^{\mathrm{move}}_{i,k}=\tau^{\mathrm{move}}_{i,k}e^{\mathrm{move}}_{k}$, where $\tau^{\mathrm{move}}_{i,k}$ denote the traveling time of the carrier vehicle when moving from its origin to the target ES $s_{i}$, which depends on the driving distance and the AgP’s vehicular speed; (ii) UAV flying cost: the energy consumption during UAV flight is $c^{\mathrm{fly}}_{i,k}=\tau^{\mathrm{fly}}_{i,k}e^{\mathrm{fly}}_{k}$, where $\tau^{\mathrm{fly}}_{i,k}$ denote the total UAV flight time spent shuttling between $s_{i}$ and its associated service region, which increases with the number and size of the tasks to be served; (iii) UAV computing cost: the energy consumed by UAV while assisting edge services is $c^{\mathrm{comp}}_{i,k}=\tau^{\mathrm{comp}}_{i,k}e^{\mathrm{comp}}_{k}=\frac{\overline{r}e_{i}^{\mathrm{E}}}{f_{k}^{\mathrm{V}}}$, where $\overline{r}$ denotes the historical average computational workload per task [^4] Since trading is conducted in the offline stage, AgPs do not know the exact sizes of the tasks they will eventually process and can only evaluate them based on historical average data. , and $\tau^{\mathrm{comp}}_{i,k}$ represent the effective computing time that the UAV dedicates to processing offloaded tasks, determined by the aggregate computational workload and the UAV’s computing capability; and (iv) Income for service provisioning: the revenue obtained from contributing edge services is denoted by $r^{\mathrm{ES}}_{i,k}$. The overall utility of $v_{k}$ can therefore be expressed as
$U^{\mathrm{V}}_{k}=\sum_{s_{i}\in\bm{\mathcal{{S}}}}x^{\mathrm{Off}}_{i,k}\omega_{1}\left(r^{\mathrm{{ES}}}_{i,k}-c^{\mathrm{move}}_{i,k}-c^{\mathrm{fly}}_{i,k}-N_{i,k}^{\mathrm{Trad}}c^{\mathrm{comp}}_{i,k}\right)+\omega_{2}c^{\mathrm{hard}},$ (2)
where $\omega_{1}$ and $\omega_{2}$ denote cost coefficients that convert energy consumption into monetary units (e.g., dollars), and $c^{\mathrm{hard}}$ represents the hardware depreciation cost.
@@ IV-A3 Utility of the Auctioneer
Since trading contracts are driven by a well-designed auction process, we involve an auctioneer (e.g., acted by a trusted third party), playing a pivotal role in coordinating interactions among players, including bids/asks collection, winning AgPs-ESs determination, and trading contract design. Accordingly, utility of the auctioneer generally depends on the difference between the total income obtained by AgPs and the expense paid by ESs:
$U^{\mathrm{Auc}}=\sum_{s_{i}\in\bm{\mathcal{S}}}\sum_{v_{k}\in\bm{\mathcal{V}}}x^{\mathrm{Off}}_{i,k}N^{\mathrm{{Trad}}}_{i,k}\left(p^{\mathrm{{ES}}}_{i,k}-r_{i,k}^{\mathrm{ES}}\right),$ (3)
IV-B Problem of Our Interest in Offline Stage
A fundamental and critical goal within an auction market is to maximize its social welfare (SW), for which we design as the collective utilities of three parties (e.g., ESs, AgPs, auctioneer), as given by:
$U^{\mathrm{SW}}=\sum_{s_{i}\in\bm{\mathcal{S}}}U^{\mathrm{ES,Off}}_{i}+\sum_{v_{k}\in\bm{\mathcal{V}}}U^{\mathrm{V}}_{k}+U^{\mathrm{Auc}}.$ (4)
@@ IV-B Problem of Our Interest in Offline Stage
s.t. $\sum_{v_{k}\in\bm{\mathcal{V}}}x^{\mathrm{Off}}_{i,k}\leq 1,\ \forall\ s_{i}\in\bm{\mathcal{S}}$ (5a)
$x^{\mathrm{Off}}_{i,k}\in\left\{0,1\right\},\forall\ s_{i}\in\bm{\mathcal{S}},\forall v_{k}\in\bm{\mathcal{V}}$ (5b)
$\sum_{s_{i}\in\bm{\mathcal{S}}}{x^{\mathrm{Off}}_{i,k}N^{\mathrm{{Trad}}}_{i,k}}\leq G_{k}^{\mathrm{V}},\forall v_{k}\in\bm{\mathcal{V}}$ (5c)
where constraint (5a) ensures that each ES is matched with a unique AgP, and (5b) enforces the binary nature of $x^{\mathrm{Off}}_{i,k}$. Constraint (5c) guarantees that the promised resources in signed contracts do not exceed the actual supply provided by $v_{k}$.
@@
IV-C Forecasting, Agent Deployment and Contracting

### v2

@@ III-B1 Modeling of SDs
We model HUs and MUs as follows:
(i) HUs: Each HU $u_{m}^{\mathrm{H}}\in\bm{\mathcal{U}}^{\mathrm{H}}$ is modeled by a 5-tuple $u_{m}^{\mathrm{H}}=\{e_{m}^{\mathrm{H,tx}},\,r_{m}^{\mathrm{H}},\,d_{m}^{\mathrm{H}},\,f_{m}^{\mathrm{H,loc}},\,\bm{C}_{m}^{\mathrm{H}}\}$, where $e_{m}^{\mathrm{H,tx}}$ is its uplink transmit power (Watts); $r_{m}^{\mathrm{H}}$ is its required computing workload (CPU cycles); $d_{m}^{\mathrm{H}}$ is its input data size (bits); $f_{m}^{\mathrm{H,loc}}$ is its local CPU frequency; and $\bm{C}_{m}^{\mathrm{H}}$ represents the set of reachable ESs within the communication coverage of $u_{m}^{\mathrm{H}}$.
(ii) MUs: Each MU $u_{n}^{\mathrm{M}}\in\bm{\mathcal{U}}^{\mathrm{M}}$ is represented by a 6-tuple $u_{n}^{\mathrm{M}}=\{\tau_{n}^{\mathrm{ddl}},\,e_{n}^{\mathrm{M,tx}},\,r_{n}^{\mathrm{M}},\,d_{n}^{\mathrm{M}},\,f_{n}^{\mathrm{M,loc}},\,\bm{C}_{n}^{\mathrm{M}}\}$, where $\tau_{n}^{\mathrm{ddl}}$ is its task deadline time; $e_{n}^{\mathrm{M,tx}}$ is its transmit power; $r_{n}^{\mathrm{M}}$ and $d_{n}^{\mathrm{M}}$ denote its workload and data size, respectively. Also, $f_{n}^{\mathrm{M,loc}}$ denotes its CPU capability, and $\bm{C}_{n}^{\mathrm{M}}$ captures SPs that can serve $u_{n}^{\mathrm{M}}$.
@@ III-B2 Modeling of SPs
We model ESs and APs as follows:
@@
(i) ESs: Each ES $s_{i}\in\bm{\mathcal{S}}$ is represented by a 6-tuple: $s_{i}=\{t_{i}^{\mathrm{E,b}},f_{i}^{\mathrm{E}},\,e_{i}^{\mathrm{E}},\,G_{i},\,l_{i}^{\mathrm{E}}\}$, where $t_{i}^{\mathrm{E,b}}$ denotes the time-slot at which $s_{i}$ begins to request resource supply to mitigate the shortage of edge resources within its own coverage area; $f_{i}^{\mathrm{E}}$ is its computing capacity (e.g., CPU cycles/s); $e_{i}^{\mathrm{E}}$ is its local power consumption (Watts); $l_{i}^{\mathrm{E}}=(x_{i}^{\mathrm{E}},y_{i}^{\mathrm{E}},h_{i}^{\mathrm{E}})$ is its 3D location. In addition, each ES $s_{i}$ is presumed to have a set of orthogonal subcarriers that can serve at most $G_{i}$ SDs simultaneously.
@@
(ii) APs: each AP $v_{k}\in\bm{\mathcal{V}}$ is modeled by a 7-tuple: $v_{k}=\{C_{k}^{\mathrm{A}},\,f_{k}^{\mathrm{A}},\,e_{k}^{\mathrm{comp}},\,e_{k}^{\mathrm{move}},\,e_{k}^{\mathrm{fly}},\,G_{k}^{\mathrm{A}},\,l_{k}^{\mathrm{A}}\}$, where $C_{k}^{\mathrm{A}}$ denotes the maximum number of UAVs carried by the vehicle; $f_{k}^{\mathrm{A}}$ captures the computing capability of each mounted UAV; $e_{k}^{\mathrm{move}}$, $e_{k}^{\mathrm{comp}}$, and $e_{k}^{\mathrm{fly}}$ denote the power consumption of ground movement, UAV computation, and UAV flight, respectively; $G_{k}^{\mathrm{A}}$ denotes the maximum number of SDs that each UAV of $v_{k}$ can serve through its limited orthogonal subcarriers; $l_{k}^{\mathrm{A}}\langle t\rangle{=}(x_{k}^{\mathrm{A}}\langle t\rangle,y_{k}^{\mathrm{A}}\langle t\rangle)$ is the ground location of the vehicle at time $t$.
III-B3 Modeling of Offline Contracts
@@
In the offline stage, a trading contract is established between each ES $s_{i}$ and AP $v_{k}$, denoted as $\mathbb{C}^{\mathrm{E}\leftrightarrow\mathrm{A}}_{i,k}=\{N^{\mathrm{Trad}}_{i,k},p_{i,k}^{\mathrm{ES}},r_{i,k}^{\mathrm{ES}}\}$, where $N^{\mathrm{Trad}}_{i,k}$ represents the quantity of traded resources, $p_{i,k}^{\mathrm{ES}}$ indicates unit payment (per task) from $s_{i}$ to contractual AP $v_{k}$, and $r_{i,k}^{\mathrm{ES}}$ refers to the unit reward (per task) of $v_{k}$ for offering services.
IV Offline Stage: Forecasting-Driven Contracting and AP Deployment
@@
We first delve into capturing the service leasing among different types of SPs, for which we systematically transform the handling of uncertain spatio-temporal service demand into proactive, incentive-compatible, and cost-aware AP deployment decisions. Our major efforts include: (i) a predictive workload forecasting module that estimates the spatiotemporal dynamics of demands across distributed ESs; (ii) a cost-aware vehicle routing strategy based on eACO, which constructs time-window-feasible candidate associations/matching of APs to ESs; (iii) a contract-design module that leverages an incentive-compatible auction mechanism to determine long-term cooperative agreements between ESs and APs, including winning trading pairs, allocated tasks, and corresponding payment rules.
@@
We let $x^{\mathrm{Off}}_{i,k}$ denote the assignment/matching between ESs and APs, where $x^{\mathrm{Off}}_{i,k}=1$ indicates that ES $s_{i}$ successfully secures the service of AP $v_{k}$ and thereby signs an offline contract, and $x^{\mathrm{Off}}_{i,k}=0$ otherwise. For notational simplicity, we define $\bm{X}^{\mathrm{Off}}=\big\{x^{\mathrm{Off}}_{i,k}\mid s_{i}\in\bm{\mathcal{S}},v_{k}\in\bm{\mathcal{V}}\big\}$ as the assignment profile.
IV-A Utilities of ESs, AP, and Auctioneer
@@ IV-A1 Utility of ESs
In the offline stage, as delivering edge services to users and leasing computing capacity can generate both costs and revenues for ESs, we model the utility of each ES $s_{i}$ as its net profit as follows:
$U^{\mathrm{ES,Off}}_{i}=\sum_{v_{k}\in\bm{\mathcal{V}}}x^{\mathrm{Off}}_{i,k}N_{i,k}^{\mathrm{Trad}}\left(\overline{p}_{i}-p^{\mathrm{ES}}_{i,k}\right),$ (1)
@@
IV-A2 Utility of APs
We model the utility of each AP as a combination of four key components: (i) Vehicular moving cost: the energy cost incurred by the vehicle traveling to ES $s_{i}$ is given by: $c^{\mathrm{move}}_{i,k}=\tau^{\mathrm{move}}_{i,k}e^{\mathrm{move}}_{k}$, where $\tau^{\mathrm{move}}_{i,k}$ denotes the traveling time of the carrier vehicle when moving from its origin to the target ES $s_{i}$, which depends on the driving distance and the AP’s vehicular speed; (ii) UAV flying cost: the energy consumption during UAV flight is given by: $c^{\mathrm{fly}}_{i,k}=\tau^{\mathrm{fly}}_{i,k}e^{\mathrm{fly}}_{k}$, where $\tau^{\mathrm{fly}}_{i,k}$ denote the total UAV flight time spent shuttling between $s_{i}$ and its associated service region, which increases with the number and size of the tasks to be served; (iii) UAV computing cost: the energy consumed by each UAV while providing resources is given by: $c^{\mathrm{comp}}_{i,k}=\tau^{\mathrm{comp}}_{i,k}e^{\mathrm{comp}}_{k}=\frac{\overline{r}e_{i}^{\mathrm{E}}}{f_{k}^{\mathrm{A}}}$, where $\overline{r}$ denotes the historical average computational workload per task [^2] Since trading is conducted in the offline stage, APs do not know the exact sizes of the tasks they will eventually process and can only evaluate them based on historical average data. , and $\tau^{\mathrm{comp}}_{i,k}$ represents the effective computing time that the UAV dedicates to processing offloaded tasks, determined by the aggregate computational workload and the UAV’s computing capability; (iv) Income for service provisioning: the revenue obtained from contributing resources is denoted by $r^{\mathrm{ES}}_{i,k}$. The overall utility of $v_{k}$ can therefore be expressed as follows:
$U^{\mathrm{A,Off}}_{k}\hskip-2.84526pt=\hskip-2.84526pt\sum_{s_{i}\in\bm{\mathcal{{S}}}}x^{\mathrm{Off}}_{i,k}\omega_{1}\left(r^{\mathrm{{ES}}}_{i,k}\hskip-2.84526pt-c^{\mathrm{move}}_{i,k}\hskip-2.84526pt-c^{\mathrm{fly}}_{i,k}-N_{i,k}^{\mathrm{Trad}}c^{\mathrm{comp}}_{i,k}\right)+\omega_{2}c^{\mathrm{hard}},$ (2)
where $\omega_{1}$ and $\omega_{2}$ denote cost coefficients that convert energy consumption into monetary units (e.g., dollars), and $c^{\mathrm{hard}}$ represents the hardware depreciation/maintenance cost.
@@ IV-A3 Utility of the Auctioneer
An auctioneer (e.g., a trusted third party), plays a pivotal role in coordinating interactions among APs and ESs, including bids/asks collection, determining winning APs-ESs associations, and trading contract design. We define the utility of auctioneer as the difference between the income obtained by APs and the expense paid by ESs:
$U^{\mathrm{Auc,Off}}=\sum_{s_{i}\in\bm{\mathcal{S}}}\sum_{v_{k}\in\bm{\mathcal{V}}}x^{\mathrm{Off}}_{i,k}N^{\mathrm{{Trad}}}_{i,k}\left(p^{\mathrm{{ES}}}_{i,k}-r_{i,k}^{\mathrm{ES}}\right).$ (3)
IV-B Problem of Interest in Offline Stage of FUSION
Our overarching goal during the offline stage of FUSION is maximizing the social welfare (SW) defined as the collective utilities of engaged parties (i.e., ESs, APs, auctioneer) as:
$U^{\mathrm{SW}}=\sum_{s_{i}\in\bm{\mathcal{S}}}U^{\mathrm{ES,Off}}_{i}+\sum_{v_{k}\in\bm{\mathcal{V}}}U^{\mathrm{A}}_{k}+U^{\mathrm{Auc}}.$ (4)
@@ IV-B Problem of Interest in Offline Stage of FUSION
s.t. $\sum_{v_{k}\in\bm{\mathcal{V}}}x^{\mathrm{Off}}_{i,k}\leq 1,~~~\forall s_{i}\in\bm{\mathcal{S}}$ (5a)
$x^{\mathrm{Off}}_{i,k}\in\left\{0,1\right\},~~~\forall s_{i}\in\bm{\mathcal{S}},~\forall v_{k}\in\bm{\mathcal{V}}$ (5b)
$\sum_{s_{i}\in\bm{\mathcal{S}}}{x^{\mathrm{Off}}_{i,k}N^{\mathrm{{Trad}}}_{i,k}}\leq G_{k}^{\mathrm{A}},~~~\forall v_{k}\in\bm{\mathcal{V}}$ (5c)
@@
IV-C Forecasting, Agent Deployment, and Contracting
We next present the offline stage that couples demand forecasting, AP deployment, and long-term contracting, with detailed interaction flow moved to Appx. B-A. In what follows, we describe the three modules, i.e., Pro-LNN-based demand forecasting, eACO-VRP-based AP deployment, and the Off-AIC2 contracting mechanism.
Unlike discrete-time recurrent neural networks (RNNs) or Transformers, LNNs [23] implement continuous-time dynamics with neuron-wise time constants and exponential-style state updates. This makes them particularly suitable for modeling time series that are sampled at irregular resource trading intervals, exhibit non-stationary spatio–temporal demand patterns, and must be processed under strict edge-computing resource constraints. In our setting, Pro-LNN takes as input the ES-side demand time series across trading rounds $\nu=1,2,\ldots$, where the time gaps $\Delta\nu$ between rounds can be irregular and some observations may be missing due to asynchronous reporting. The predictor therefore must stay stable under such irregular $\Delta\nu$ while capturing both short-term and long-term trends, motivating our LNN-based architecture.
Let $N_{i}^{\mathrm{Dem},(\nu)}$ denote the task demand of ES $s_{i}$ at the $\nu$-th trading round. We form an input vector $\mathbf{z}_{i}^{(\nu)}=\big[N_{i}^{\mathrm{Dem},(\nu)},\mathbf{x}_{i}^{(\nu)}\big]$, where $\mathbf{x}_{i}^{(\nu)}$ stacks optional exogenous features at round $\nu$, such as time-of-day and day-of-week indicators, ES-type and capacity descriptors, and local mobility/load statistics extracted from recent logs. Given a length-$L$ history of past observations $\{\mathbf{z}_{i}^{(\nu-L+1)},\ldots,\mathbf{z}_{i}^{(\nu)}\}$, Pro-LNN encodes this sequence via a liquid cell with neuron-wise time constants, integrates the measurements into a hidden state $\mathbf{h}_{i}^{(\nu)}$, and then applies a shallow readout network to produce an $H$-step forecast $\{\hat{N}_{i}^{\mathrm{Dem},(\nu+1)},\ldots,\hat{N}_{i}^{\mathrm{Dem},(\nu+H)}\}$. In a nutshell, Pro-LNN is trained with a composite loss, which is an $H$-step MSE between the forecasts and the ground-truth demands, and we additionally use a multi-step consistency regularizer that rolls the predictor forward in closed-loop over $H$ steps and penalizes the accumulated prediction errors at future trading rounds. These forecasts are computed offline before each trading round and then fed into the eACO-VRP module for ES recommendation and AP route pre-planning, as well as the auction module for ES–AP contract design. The detailed update equations and training pseudocode of our Pro-LNN are deferred to Appx. B-B to avoid detracting from the main discussion.
Within the proposed offline stage of FUSION, once Off-AIC2 has identified a feasible trading region and candidate ES–AP pairs, the auction platform still needs to decide, for each winning AP, which ESs should be physically visited and in what order, under time-window and capacity constraints. This induces a capacitated vehicle-routing problem with time windows and heterogeneous travel costs and utilities. Since our APs (e.g., UAVs or service vehicles) and the auction platform operate under tight energy and latency budgets and should periodically recompute routes as demand forecasts change, we adopt a lightweight ant colony optimization (ACO) paradigm and tailor it into an eACO-VRP that can be invoked as a subroutine inside Off-AIC2, balancing solution quality with low computational complexity and energy cost.